Can you help me please,
Find all values of u satisfying the equation
7/u-1 =-6 + 2/u+2
.
(If there is more than one solution, separate them with commas





Answer by checkley77(12844)

(7/u)-1 =-6 + (2/u)+2
7/u-2/u=-6+2+1
(7-2)/u=-3
5/u=-3
-3u=5
u=5/-3
u=-5/3 answer.
OR:
7/(u-1)=-6+2/(u+2)
7/(u-1)-2/(u+2)=-6
[7(u+2)-2(u-1)]/(u-1)(u+2)=-6
(7u+14-2u+2)/(u^2+u-2)=-6
(5u+16)/(u^2+u+2)=-6
-6(u^2+u-2)=5u+16
-6u^2-6u+12=5u+16
-6u^2-6u-5u+12-16=0
-6u^2-11u-4=0
6u^2+11u+4=0
(3u+4)(2u+1)=0
3u+4=0
3u=-4
u=-4/3 answer.
2u+1=0
2u=-1
u=-1/2 answer.
(-4/3,-1/2)